For a sum of money invested at

4.44.4​%

compounded

semi dash annuallysemi-annually

for

33

years state the following values.

​(a) the number of compounding periods

​(b) the periodic rate of interest

​(c) the compounding factor left parenthesis 1 plus i right parenthesis Superscript n(1+i)n

​(d) the numerical value of the compounding factor
Answer #1

a). Number of compounding periods = Number of Years * Compounding periods in a year

= 33 * 2 = 66

b). Periodic Interest Rate = Annual Rate / Compounding periods in a year = 4.4% / 2 = 2.2%

c). Compounding Factor = (1 + i)n = (1 + 0.022)66

d). Numerical Value of Compounding factor = (1 + 0.022)66 = 4.2049